OpenStudy (chelseatheruler):

Given the graph y=x^2+4x+3. How to find the vertex of the quadratic equation.

OpenStudy (jdoe0001):

\(\bf \textit{vertex of a parabola}\\ \quad \\ y = {\color{red}{ a}}x^2+{\color{blue}{ b}}x+{\color{green}{ c}}\qquad \qquad \quad \left(-\cfrac{{\color{blue}{ b}}}{2{\color{red}{ a}}}\quad ,\quad {\color{green}{ c}}-\cfrac{{\color{blue}{ b}}^2}{4{\color{red}{ a}}}\right)\)

OpenStudy (chelseatheruler):

But, how do I solve this problem?

OpenStudy (chelseatheruler):

I don't understand how.

OpenStudy (jdoe0001):

you plug in your values \(\bf \textit{vertex of a parabola}\\ \quad \\ y = {\color{red}{ 1}}x^2{\color{blue}{ +4}}x{\color{green}{ +3}}\qquad \qquad \quad \left(-\cfrac{{\color{blue}{ b}}}{2{\color{red}{ a}}}\quad ,\quad {\color{green}{ c}}-\cfrac{{\color{blue}{ b}}^2}{4{\color{red}{ a}}}\right)\)

OpenStudy (chelseatheruler):

So it would be 4/2 and 3-4/4?

OpenStudy (chelseatheruler):

Which would be (2,-3)?

OpenStudy (jdoe0001):

close, but not there yet \(\bf \left(-\cfrac{{\color{blue}{ 4}}}{2{\color{red}{ (1)}}}\quad ,\quad {\color{green}{ 3}}-\cfrac{{\color{blue}{ 4}}^2}{4{\color{red}{ (1)}}}\right)\)